OpenStudy (anonymous):

Williamston Daily is a new daily newspaper. When they first started printing three years ago they had 5600 home delivery subscribers. The number of subscribers increases by 71 every month. Write a linear model that represents the number of subscribers, B, for the number of months, n, the paper has been printed.

OpenStudy (anonymous):

answer choices a. B = 71n + 5600 c. B = 213n + 5600 b. B = 71n + 3 d. B = 5600n + 71

Mercury:

Asked & answered here: https://questioncove.com/study#/updates/5f8e82103f2de8644ebeb200